OpenStudy (anonymous):

A company distributes free crayon boxes to all the students of x schools. Each school has (x + 5) classes. The number of students in each class is 2 less than the number of classes in each school. Each student is given two boxes of crayons Part 1: Write an expression to show the total number of crayon boxes distributed by the company in x schools.

OpenStudy (ranga):

Number of schools = x Number of classes in each school = (x + 5) Number of students in each class = Number of classes in each school - 2 = x + 5 - 2 = (x + 3) Number of crayons for each student = 2 Can you take it from here?

OpenStudy (ranga):

Total number of crayons distributed = Number of crayons for each student * Number of students in each class * Number of classes in each school * Number of schools

OpenStudy (anonymous):

This was very helpful, and here's what it'd look like: \[A = l • w \] \[A = (6x^2 + 3x - 2) • (x^3 - 2x + 5)\]

OpenStudy (anonymous):

Solve from there, once you know what to do my homie.